# Spatio-temporal Fourier Transformer (StFT) for Long-term Dynamics Prediction

### v1

@@ 3.1 StFT Block
Figure 2 illustrates the architecture of the StFT block. Assuming the collected data samples $\tilde{u}_{t}$ are mapped to a tensor of $W\times H\times C$, where $W,H$ are spatial dimensions, and $C$ is for the number of physical variables. Each StFT block is designed to handle a specific scale; therefore, by employing a specific patch size, we partition the input at a corresponding level of granularity. To ensure that the first block captures the coarsest features or the largest scale, we set the patch size $P_{w}^{1}\times P_{h}^{1}$ to a large value, allowing it to model broad spatial structures effectively. Overlaps between patches are not considered in this section. As a result, $\frac{W}{P_{w}^{1}}\times\frac{H}{P_{H}^{1}}\times C$ patches or tokens are fed into the variable mixer block, where tokens corresponding to different physical processes but sharing the same spatial and temporal locations are mixed into a single token. Following this, two transformation paths are performed: one in the spatio-temporal domain, and the other in the frequency domain. In the spatio-temporal path, the tokens first pass through a spatio-temporal embedder, after which the embeddings are processed by multiple standard transformer layers. In the frequency path, the tokens are first processed by a 2D/3D Fourier transform, where only low-frequency components are retained. These low-frequency components are then passed through a frequency embedder to obtain frequency embeddings. Subsequently, these frequency embeddings are fed to the standard transformer layers for further processing. Finally, an inverse 2D/3D Fourier transform is applied to map the frequency tokens back to the spatio-temporal domain. Once the tokens from both paths are obtained, a linear layer is used to project them back to the original space, where each token has a shape of $P_{w}^{1}\times P_{h}^{1}$. Next, a detokenizer yields the first block prediction $u_{t+1}^{1}$. Note that the number of tokens remains unchanged within the transformer block. Each patch represents a significant portion of the historical snapshots, encapsulating macroscopic structural features. This coarse-level partitioning reduces the complexity of modeling fine-grained details. By maintaining a lower granularity, the model prioritizes structural coherence over extraneous details, enabling it to focus on capturing and predicting global relationships between regions more effectively.
@@ 3.2 A Hierarchy of StFT Blocks
In the subsequent blocks, we shift our focus to smaller scales with fine details. Consequently, we concatenate the prediction $u_{t+1}^{1}$ with the input $\tilde{u}_{t}$, and consider this combined input as the input for the next StFT block. We further subdivide each patch from the previous level into smaller patches. The smaller patches allows for less information to be aggregated within a single patch, thereby minimizing the risk of losing local variations and enhancing the richness and informativeness of the fine-scale representation. As a result, the subdivision increases the resolution or the total number of patches to the next StFT block. By leveraging the finer granularity of the patches to focus on smaller regions, it allows the model to better localize features and capture their details. In addition, conditioning on the coarser prediction allows the models to iteratively refine its estimates, beginning with a broad global summary of the corresponding regions. This iterative refinement process mirrors how humans process images, first focusing on global structures and then honing in on finer details. Through the repeated subdivision of the patches, the model progressively refines its predictions across multiple scales.
Figure 1: Overview of the proposed StFT model. Top right: the model predicts $u_{t+1}$ using the past $k$ snapshots $\tilde{u}_{t}$, employing $L$ spatial refinements from coarse to fine scales through the proposed spatiotemporal Fourier transformer blocks. Bottom left: an illustration of the overlapping tokenizer, where the patch size is $2\times 2$, and the overlapping number is $1\times 1$.
Figure 2: Illustration of the proposed spatiotemporal Fourier transformer block for StFT model: first, the past snapshots $\tilde{u}_{t}$ and a coarser prediction from the previous layer are temporally stacked. The stacked discretized function is then passed to the overlapping tokenizer to generate tokens for each variable. Next, tokens corresponding to different variables in the same spatial and temporal locations are mixed through a variable mixer. Following this, two paths of transformation are applied on the tokens. After passing through the overlapping detokenizer, a finer prediction for timestamp $t+1$ is obtained.
3.3 Residual Refinement Based on Flow Matching
Finally, the model refines its predictions through a rectified flow block, which belongs to the family of flow matching models (Liu, 2022; Liu et al., 2022a). Flow matching is formulated as an ordinary differential equation on time $t\in[0,1]$,
$\frac{d}{dt}\psi_{t}(x)=\nu_{t}(\psi_{t}(x)),$ (3)
@@ 3.3 Residual Refinement Based on Flow Matching
where the learnable velocity field $\nu_{t}$ directs the transformation of each sample $X_{0}$ from a source distribution $p_{0}$, typically a Gaussian distribution, toward the target distribution $X_{1}\sim p_{1}$, with $p_{1}$ representing the data distribution. If we prescribe the velocity field $\nu_{t}$ such that it guides every sample along a straight-line trajectory from $X_{0}$ to $X_{1}$, it is referred to as a rectified flow.. In this case, $X_{t}$ represents the linear interpolation across the entire timespan between $X_{0}$ and $X_{1}$, which can be expressed as
$X_{t}=tX_{1}+(1-t)X_{0}.$ (4)
We employ a parameterized $\mathcal{M}_{\theta_{g}}$ to approximate $\nu_{t}$, leading to the following learning objective:
$\mathcal{L}(\theta_{g})=\mathbb{E}_{t,X_{0},X_{1}}\left\|\mathcal{M}_{\theta_{g}}(X_{t})-(X_{1}-X_{0})\right\|^{2}.$ (5)
In our model, the block takes the prediction $u_{t+1}^{L}$ from the final block of StFT and the past snapshots $\tilde{u}_{t}$ as conditioning inputs. Its objective is to generate the distribution of residuals $y-\sum_{j}u_{t+1}^{j}$, where $y$ is the ground truth for the solution at $t+1$. Our training loss then becomes:
$\mathbb{E}_{X_{0}\sim\mathbf{N}(0,\mathbf{I}),t\sim(0,1)}[(\mathcal{M}_{\theta_{g}}(\tilde{u}_{T},\mathcal{F}_{\theta_{d}}(\tilde{u}_{T}),(1-t)*X_{0}$
$+t*(y-\sum_{j}u_{T+1}^{j}))-(y-\sum_{j}u_{T+1}^{j}-X_{0}))^{2}],$ (6)
where we use $T$ to denote the timestamp in $\tilde{u}_{T}$ and $u_{T+1}$.
Algorithm 1 $\mathrm{StFT}$
  Initialize: For each StFT block $l\in[1,L]$, patch sizes $(p_{h},p_{w})_{l}$, truncation frequencies $(m_{h},m_{w})_{l}$, overlapping numbers $(o_{h},o_{w})_{l}$, $u_{t+1}^{0}$ as $\mathsf{None}$, and output $u_{t+1}$ as $\mathsf{zeros}$.
  Input: $\tilde{u}_{t}\leftarrow(u_{t},u_{t-1},\dots,u_{t-k+1})$
  Output: $u_{t+1}$
[…]
  $v\leftarrow v\text{-th variable in }V$
  for $l=1$ to $L$ do
     $x_{t}\leftarrow\mathsf{TemporalStacking}(\tilde{u}_{t},\,u_{t+1}^{l-1})$
     $\{x_{i,t}\}_{v}\leftarrow\mathsf{OLTokenizer}(x_{t},p_{h},p_{w},o_{h},o_{w})$
     Frequency Path
       $\bar{x}_{i,t,1}\leftarrow\mathsf{VariableMixer}^{1}(\{x_{i,t}\}_{v\in V})$
       $f_{i,t}\leftarrow\mathsf{FreqEmbedder}(\mathsf{FFTFilter}(\bar{x}_{i,t,1},m_{h},m_{w}))$
       $f_{i,t}\leftarrow\mathsf{TransformerBlock}^{1}(f_{i,t})$
       $\{x_{i,t+1,1}\}_{v}\leftarrow\mathsf{Linear}^{1}(\mathsf{iFFT}(f_{i,t},m_{h},m_{w}))$
     Spatiotemporal Path
       $\bar{x}_{i,t,2}\leftarrow\mathsf{VariableMixer}^{2}(\{x_{i,t}\}_{v\in V})$
       $e_{i,t}\leftarrow\mathsf{StEmbedder}(\bar{x}_{i,t,2})$
       $e_{i,t}\leftarrow\mathsf{TransformerBlock}^{2}(e_{i,t})$
       $\{x_{i,t+1,2}\}_{v}\leftarrow\mathsf{Linear}^{2}(e_{i,t})$
     Merge
       $\{x_{i,t+1}\}_{v}\leftarrow\{x_{i,t+1,1}\}_{v}+\{x_{i,t+1,2}\}_{v}$
     $u_{t+1}^{l}\leftarrow\mathsf{OLDetokenizer}(\{x_{i,t+1}\}_{v},p_{h},p_{w},o_{h},o_{w})$
     $u_{t+1}\leftarrow u_{t+1}+u_{t+1}^{l}$
  end for
  return $u_{t+1}$
3.4 Overlapping Tokenizer and Detokenizer
Next, we introduce the overlapping tokenizer and the overlapping detokenizer in our model. The overlapping tokenizer is illustrated in bottom left of Figure 1. For any adjacent two patches, they share a boundary. For instance, a $3\times 3$ input generates four $2\times 2$ patches, with gray boxes representing the shared boundaries. The overlapping detokenizer reconstructs these shared boundaries by computing the mean over the overlapping regions of any two adjacent patches. This strategy effectively mitigates discontinuity issues at the patch boundaries, which is particularly important when the target function that the model aims to predict is continuous and smooth. In addition, it improves feature learning and extraction by incorporating shared boundary features into the patch embeddings. This not only preserves local continuity, but also enables the model to capture spatial relationships between patches more effectively, particularly during the self-attention mechanism.
@@ 4.1 Datasets
In this section, we consider three spatio-temporal multi-physics systems arising from time-dependent PDEs of a variety of complexities, including a high-dimensional plasma dynamics system based on reconstructed equilibrium of DIII-D experimental discharges, a 2D incompressible Navier-Stokes equation in velocity-pressure form within a squared boundaries driven by an external force, and a viscous shallow-water equation modeling the dynamics of large-scale atmospheric flows on a spherical domain. The problem setup and data generation are detailed in Appendix A.
4.2 Experimental Setup and Baselines
@@
Our goal is to simulate long-time trajectories given a few initial observations. This task is particularly challenging due to the multiple correlated variables present in the Navier-Stokes and plasma magnetohydrodynamics (MHD), with the test trajectories consisting of snapshots that vary from $71$ to $244$. We employ an autoregressive framework for all the methods. During training, each model utilizes five historical snapshots to predict the subsequent snapshot. For testing, given the initial five snapshots of a trajectory, all models forecast the entire trajectories in an autoregressive manner, iteratively generating future states based on the previously forecasted outputs from the trained model.
Figure 3: Temporal evolution of perturbed electrostatic potential $\delta\phi$ contours predicted by different models: StFT-F, StFT, FNO, ViT and U-Net. Significant phase differences between the predictions of the models appear after $T=59$, where StFT and StFT-F perform stable across the forecasting window.
@@
We evaluate these datasets using the following well-known and state-of-the-art methods for comparison: autoregressive Fourier Neural Operator (AR-FNO), autoregressive U-Net (AR-UNet), and autoregressive vision transformer (AR-ViT). For FNO, we use the authors’ open source implementation. For U-Net, we employ the implementation of the modified U-Net as evaluated in the recent BubbleML work (Hassan et al., 2023), where the modified U-Net demonstrated superior performance across most of their proposed benchmarks.
@@
We divide the trajectories of each dataset into training, test, and validation sets. For each method, we identify the tunable hyperparameters, specify a range for each hyperparameter, and conduct a grid search across all the hyperparameters. For our methods, we implement both StFT, the deterministic component of our model, and StFT-F, our model with a generative residual refinement block. First, we evaluate the forecasting performance of all methods by calculating the mean $L_{2}$ relative error over the test trajectories. For StFT-F, in order to obtain the mean prediction for all the test trajectories, we generate $50$ prediction samples at each autoregressive step, and then inject their mean into the next step. Additionally, we assess the uncertainty quantification capability of our method by sampling $100$ trajectories for each test case. For each trajectory, at each autoregressive step, we generate a single prediction sample, which is subsequently fed into the next autoregressive step to iteratively foreacst the full sequence. We ensure that all models are thoroughly trained by running sufficient epochs. We used the AdamW optimizer with a learning rate of $1e^{-4}$ to train those models on an A100 GPU with a total memory capacity of 40 GB. A comprehensive list of all the hyperparameters along with their respective ranges used for grid search is provided in the Appendix C.

### v2

@@ 3.1 StFT Block
Overlapping tokenizer. Each StFT block first tokenizes the discretized functions $\tilde{u}_{t}$ of shape $T\times W\times H\times C$, where $T$ is the temporal dimension, $W,H$ are the spatial dimensions, and $C$ denotes the number of physical variables. We apply tokenization along the spatial dimensions for each variable. To enhance spatial continuity while minimizing visual artifacts, we propose to use an overlapping tokenizer (OLT) and detokenizer (OLDT) that allows adjacent patches to share boundaries through overlapping regions. For instance, as shown in Figure 1, a $3\times 3$ input generates four $2\times 2$ patches with a $1\times 1$ overlap, where the overlapping areas (indicated in gray) are shared between patches. During detokenization, overlapping regions are reconstructed by averaging the corresponding values from neighboring patches. This strategy effectively mitigates discontinuity issues at the patch boundaries, which is particularly important for accurately representing smooth and continuous target functions. Furthermore, incorporating shared boundaries into patch embeddings enriches feature representation and extraction from fine-scale structures.
Variable mixer. Each StFT block is designed to handle a specific scale; therefore, by employing a specific patch size, we partition the input at a corresponding level of granularity. To ensure that the first block captures the coarsest features or the largest scale, we set the patch size $p_{w}^{1}\times p_{h}^{1}$ to a large value, allowing it to model broad spatial structures effectively. As a result, $\mathcal{O}(\frac{W}{p_{w}^{1}}\times\frac{H}{P_{h}^{1}}\times C)$ patches are fed into the variable mixer, where patches corresponding to different physical processes but sharing the same spatial domain are mixed into a single token. Following this step, two transformation paths are performed: one is in the spatio-temporal domain, which operates self-attention on spatio-temporal embeddings; the other is in the frequency domain, which operates on frequency embeddings.
Frequency embeddings. The tokens are first processed by a 2D/3D Fourier transform, where only low-frequency components are retained. These low-frequency components are then passed through a frequency embedder to obtain frequency embeddings $f_{t}$. Subsequently, these frequency embeddings are fed to the standard transformer layers for mixing information and nonlinear transformation in the frequency domain. Finally, an inverse 2D/3D Fourier transform and a linear projection are applied to map the frequency embeddings back to the spatio-temporal domain.
Spatio-temporal embeddings. The same set of tokens first pass through a spatio-temporal embedder, after which the spatio-temporal embeddings $e_{t}$ are processed by multiple standard transformer layers for mixing correlations and nonlinear transformation. Finally a separate linear projection is applied to get predictions for each patch of $u_{t+1}$.
Next, an overlapping detokenizer yields the first block prediction $u_{t+1}^{1}$. Each token represents a significant portion of the historical snapshots, encapsulating macroscopic structural features. This coarse-level partitioning reduces the complexity of modeling fine-grained details. By maintaining a lower granularity, the model prioritizes structural coherence over extraneous details, enabling it to focus on capturing and predicting global relationships between regions more effectively.
@@ 3.2 A Hierarchy of StFT Blocks
Figure 2: A patch is subdivided into smaller patches for a hierarchical learning.
@@
In the subsequent StFT blocks, we shift our focus to smaller scales with fine details. Consequently, we concatenate the prediction $u_{t+1}^{1}$ with the input $\tilde{u}_{t}$, and consider this combination as the input for the next StFT block. We further subdivide each patch from the previous StFT block into smaller patches. The smaller patches allow for less information to be aggregated within a single patch, thereby minimizing the risk of losing local variations and enhancing the richness and informativeness of the fine-scale representation. By leveraging the finer granularity of the patches to focus on smaller regions, it allows the model to better localize features and capture their details. In addition, conditioning on the coarser prediction allows the models to iteratively refine its estimates, beginning with a broad global summary of the corresponding regions. Through the repeated subdivision of the patches, the model progressively refines its predictions across multiple scales. As shown in Figure 1, the StFT model predicts $u_{t+1}$ and applies a total of $L$ multi-scale spatial refinements.
3.3 Residual Refinement and Uncertainty Estimation Based on Flow Matching
@@
Finally, the model refines its deterministic predictions through a rectified flow block, which belongs to the family of flow matching models (liu2022rectified, ; liu2022flow, ). Flow matching is formulated as an ordinary differential equation in time $\tau\in[0,1]$, $\frac{d}{d\tau}\psi_{\tau}(x)=\nu_{\tau}(\psi_{\tau}(x))$, where the learnable velocity field $\nu_{\tau}$ directs the transformation of each sample $X_{0}$ from a source distribution $p_{0}$, typically a Gaussian distribution, toward the target distribution $X_{1}\sim p_{1}$ with $p_{1}$ representing the data distribution. If we prescribe the velocity field $\nu_{\tau}$ such that it guides every sample along a straight-line trajectory from $X_{0}$ to $X_{1}$, it is referred to as a rectified flow. In this case, $X_{\tau}$ represents the linear interpolation across the entire timespan between $X_{0}$ and $X_{1}$, which can be expressed as $X_{\tau}=\tau X_{1}+(1-\tau)X_{0}$. We employ a parameterized $\mathcal{M}_{\theta_{g}}$ to approximate $\nu_{\tau}$, leading to the following learning objective: $\mathcal{L}(\theta_{g})=\mathbb{E}_{\tau,X_{0},X_{1}}\left\|\mathcal{M}_{\theta_{g}}(X_{\tau},\tau)-(X_{1}-X_{0})\right\|^{2}$. In our model, the rectified flow block takes the deterministic prediction $u_{t+1}$ from the composition of $L$ StFT blocks and the observations $\tilde{u}_{t}$ as conditioning inputs. Its objective is to generate the distribution of residuals $r_{t+1}=y-\sum_{j}u_{t+1}^{j}$, where $y$ is the ground truth for the solution at $t+1$. Our training loss then becomes:
$\mathbb{E}_{X_{0}\sim\mathbf{N}(0,\mathbf{I}),\tau\sim(0,1)}[(\mathcal{M}_{\theta_{g}}(\tilde{u}_{t},\mathcal{F}_{\theta_{d}}(\tilde{u}_{t}),\tau,X_{\tau})-(r_{t+1}-X_{0}))^{2}],$ (2)
where $X_{\tau}$ is the linear interpolation between the source sample $X_{0}$ and the target $r_{t+1}$.
1Inputs: history $\tilde{u}_{t}=(u_{t},\dots,u_{t-k+1})$
2 Initialize: blocks $l\in[1,L]$, patch sizes $p_{h_{l},w_{l}}$, truncation modes $m_{h_{l},w_{l}}$, overlaps $o_{h_{l},w_{l}}$, $u_{t+1}^{0}$ as $\mathrm{None}$, $u_{t+1}$ as $\mathbf{0}$
3 $v\!\leftarrow\!\text{var idx}$
4 for $l=1,\dots,L$ do
5  $x_{t}\leftarrow\mathrm{TemporalStacking}(\tilde{u}_{t},\,u_{t+1}^{l-1})$;
6  $\{x_{t}\}_{v}\leftarrow\mathrm{OLT}(x_{t},\,p_{h,w},\,o_{h,w})$;
7  $\{x_{t+1}\}_{v}:$ Invoke Freq. & S-t. paths
8  $u_{t+1}^{l}\leftarrow\mathrm{OLDT}(\{x_{t+1}\}_{v},\,p_{h,w},\,o_{h,w})$
9  $u_{t+1}\leftarrow u_{t+1}+u_{t+1}^{l}$
10
Return: $u_{t+1}$
Algorithm 1  StFT
1Frequency Path: $\bar{x}_{t,1}\!\leftarrow\!\mathrm{VariableMixer}^{1}(\{x_{t}\}_{v})$ $f_{t}\!\leftarrow\!\mathrm{FFTFilter}(\bar{x}_{t,1},m_{h},m_{w})$ $f_{i,t}\!\leftarrow\!\mathrm{FreqEmbedder}(f_{i,t})$ $f_{t}\!\leftarrow\!\mathrm{TransformerBlock}^{1}(f_{t})$ $\{x_{t+1,1}\}_{v}\!\leftarrow\!\mathrm{iFFT}(f_{t},m_{h},m_{w})$ $\{x_{t+1,1}\}_{v}\!\leftarrow\!\mathrm{Linear}^{1}(\{x_{t+1,1}\}_{v})$
2Spatiotemporal Path: $\bar{x}_{t,2}\!\leftarrow\!\mathrm{VariableMixer}^{2}(\{x_{t}\}_{v})$ $e_{t}\!\leftarrow\!\mathrm{StEmbedder}(\bar{x}_{t,2})$ $e_{t}\!\leftarrow\!\mathrm{TransformerBlock}^{2}(e_{t})$ $\{x_{t+1,2}\}_{v}\!\leftarrow\!\mathrm{Linear}^{2}(e_{t})$
Merge: $\{x_{t+1}\}_{v}\leftarrow\{x_{t+1,1}\}_{v}+\{x_{t+1,2}\}_{v}$
Algorithm 2  StFT block: Freq. and S-t. Paths
@@ 4.1 Datasets
In this section, we consider three spatio-temporal multi-physics systems arising from time-dependent PDEs of a variety of complexities, including a high-dimensional plasma dynamics system based on reconstructed equilibrium of DIII-D experimental discharges (bai2025transfer, ), a 2D incompressible Navier-Stokes equation in velocity-pressure form within a square domain driven by an external force, and a viscous shallow-water equation modeling the dynamics of large-scale atmospheric flows on a spherical domain. The problem setup and data generation are detailed in Appendix A.
4.2 Experimental Setup
@@
Long-term multi-physics prediction up to a horizon of 244 timesteps. Our goal is to simulate long-time trajectories given a few initial observations. This task is particularly challenging due to the multiple correlated variables present in the Navier-Stokes and plasma magnetohydrodynamics (MHD), with the test trajectories consisting of snapshots that vary from $71$ to $244$. We employ an autoregressive framework for all the methods: during training, each model utilizes five historical snapshots to predict the next one in a forward pass. At test time, given the initial five snapshots of a trajectory, all models autoregressively generate the entire trajectories by iteratively predicting future states based on their own previous outputs.
Figure 3: Results of autoregressive prediction in $L_{2}$ relative error (log scale) across the timespan: (left) perturbed parallel vector potential $\delta A_{\parallel}$ in plasma MHD, and (right) magnitude of velocity in shallow-water equation. The shaded region indicates the uncertainty distribution of $\sigma$ in the relative error of StFT-F. For a given error threshold, StFT-F maintains accuracy over at least twice the time horizon compared to the baselines.
Table 1: Quantitative results of our model and baselines in the same autoregressive framework: relative $L_{2}$ error over three spatiotemporal prediction systems. AR-StFT refers to the deterministic model’s results. AR-StFT-F denotes the probabilistic model with residual refinement of StFT. All models have been subjected to hyperparameter tuning to ensure fair and optimal performance comparisons.
Dataset Variable(s) AR-StFT AR-StFT-F AR-DPOT AR-FNO AR-ViT AR-UNetb
Plasma MHD $\delta\phi$ 2.80e-2 2.24e-2 1.04e-1 2.28e-1 1.73e-1 1.02e0
$\delta A_{\parallel}$ 2.45e-2 2.30e-2 8.36e-2 2.30e-1 3.24e-1 8.13e-1
$\delta B_{\parallel}$ 3.05e-2 2.66e-2 8.98e-2 2.33e-1 1.95e-1 7.79e-1
$\delta n_{e}$ 2.84e-2 2.45e-2 8.64e-2 2.33e-1 2.08e-1 1.01e0
$\delta n_{i}$ 3.28e-2 2.93e-2 8.76e-2 2.33e-1 2.18e-1 1.04e0
$\delta u_{e}$ 3.99e-2 3.73e-2 9.59e-2 3.18e-1 2.99e-1 6.96e-1
Navier-Stokes $u$ 3.38e-2 3.30e-2 4.67e-2 4.46e-2 5.09e-2 6.16e-2
$v$ 3.60e-2 3.17e-2 4.52e-2 4.57e-2 4.60e-2 6.15e-2
pp 5.16e-2 4.44e-2 6.18e-2 5.90e-2 7.03e-2 7.84e-2
Shallow-Water $\mathbf{V}$ 6.25e-2 6.53e-2 7.97e-2 9.53e-2 1.33e-1 2.02e-1
@@
Baseline setup. We evaluate these datasets using the following well-known and state-of-the-art methods for comparison: autoregressive AFNO, autoregressive Fourier Neural Operator (AR-FNO), autoregressive U-Net (AR-UNet), and autoregressive vision transformer (AR-ViT). More specifically, for AFNO, we used the latest variant from the DPOT work  (hao2024dpot, ), which is enhanced with a temporal aggregation layer and improved expressivity through the removal of enforced sparsity. For FNO, we use the authors’ open-source implementation. For U-Net, we employ the implementation of the modified U-Net as evaluated in the recent BubbleML work (hassan2023bubbleml, ), where the modified U-Net was initially used in PDEBench work  (takamoto2022pdebench, ), and demonstrated superior performance over their baselines.
@@
Validation and hyperparameter tuning. We divide the trajectories of each dataset into training, test, and validation sets. For each method, we identify the tunable hyperparameters, specify a range for each hyperparameter, and conduct a grid search. For our methods, we implement both StFT, the deterministic component of our model, and StFT-F, our model with a generative residual refinement block. We use the AdamW optimizer with a learning rate of $1e^{-4}$ to train those models on an A100 GPU. We ensure that all models are fairly and thoroughly trained by imposing an identical time budget across all models. More specifically, we impose 24 GPU hours on the plasma MHD and Navier-Stokes datasets, and 48 GPU hours on the shallow-water dataset. A comprehensive list of all the hyperparameters along with their respective ranges is provided in the Appendix E.
Evaluation metrics. First, we evaluate the forecasting performance by calculating the mean $L_{2}$ relative error. For StFT-F, in order to obtain the mean prediction, we generate $50$ stochastic predictions at each autoregressive step, and then inject their mean into the next step. Additionally, we assess the uncertainty quantification capability of StFT-F by sampling $100$ trajectories for each test case. For each trajectory, at each autoregressive step, we generate a single prediction, which is subsequently fed into the next autoregressive step to iteratively forecast the full sequence.